# Lifelong Knowledge Editing for Vision Language Models with Low-Rank Mixture-of-Experts

### v1

@@ Abstract
Model editing aims to correct inaccurate knowledge, update outdated information, and incorporate new data into Large Language Models (LLMs) without the need for retraining. This task poses challenges in lifelong scenarios where edits must be continuously applied for real-world applications. While some editors demonstrate strong robustness for lifelong editing in pure LLMs, Vision LLMs (VLLMs), which incorporate an additional vision modality, are not directly adaptable to existing LLM editors. In this paper, we propose LiveEdit, a Lifelong vision language model Edit to bridge the gap between lifelong LLM editing and VLLMs. We begin by training an editing expert generator to independently produce low-rank experts for each editing instance, with the goal of correcting the relevant responses of the VLLM. A hard filtering mechanism is developed to utilize visual semantic knowledge, thereby coarsely eliminating visually irrelevant experts for input queries during the inference stage of the post-edited model. Finally, to integrate visually relevant experts, we introduce a soft routing mechanism based on textual semantic relevance to achieve multi-expert fusion. For evaluation, we establish a benchmark for lifelong VLLM editing. Extensive experiments demonstrate that LiveEdit offers significant advantages in lifelong VLLM editing scenarios. Further experiments validate the rationality and effectiveness of each module design in LiveEdit.  [^1] Source codes are provided in the supplementary material and will be released upon acceptance.
@@ 1 Introduction
Generation of Low-Rank Experts: Mixture-of-Experts (MoE) is a technique used to solve complex tasks by combining multiple “experts”, with each expert being responsible for specific data patterns or sub-tasks [26, 27]. In the context of VLLM editing, we treat VLLM’s adherence to a single editing sample as a sub-task. Thus, in lifelong editing, each new incoming editing sample corresponds to a specific low-rank expert, which is responsible for adjusting the model’s response accordingly. Typically, each component in MoE requires training on sub-tasks to obtain its weights. However, directly fitting on editing samples results in poor generalization. Furthermore, performing separate training for each new editing sample is inefficient. Based on the above analysis, we construct an expert generator to generate low-rank experts for each new editing sample. To ensure that the generated experts can effectively refine the VLLM response, we train the expert generator to align the VLLM, with the relevant editing metrics: reliability, textual and modal generality, and textual and modal locality [22]. We store the generated experts in an expert repository. For a new input instance, LiveEdit will select relevant experts from the repository and merge their adapted responses using hard and soft routing, as detailed below.
Hard and Soft Routing: To route and utilize these experts during inference, we propose a two-stage routing strategy. In the first phase, we extract visual features relevant to the edit sample by interacting between the prompt and the visual representation. For each incoming sample, we do the same visual feature extraction and compare it with those of the edit samples to filter out visually irrelevant experts, thereby routing the input instance only to visually relevant experts. This hard routing effectively removes irrelevant visual and textual noise. However, it does not take into account the specific semantics within the edit text, and this coarse-grained routing may lead to the selection of multiple visually matched but text-irrelevant editing samples. In the second stage, to suppress irrelevant samples, we implement multi-expert fusion, or soft routing, by computing the similarity between the input query and the edit text applied in the first stage. By combining absolute and relative weights, relevant and irrelevant editing samples are assigned higher and lower weights, respectively.
@@ 2.2 Model Editing
Model Editing for LLMs: We offer a concise summary of techniques for editing LLMs. (1) Locate-Then-Edit: This method first pinpoints the parameters connected to particular knowledge, then directly modifies these targeted parameters [40]. ROME [17] uses causal mediation analysis to pinpoint the section requiring editing. MEMIT [18] and WILKE [21], building on ROME, facilitate the multiple knowledge editing. (2) Meta Learning: This approach uses a hyper-network to derive the updated weights necessary for editing [41, 42, 43, 44]. (3) Additional Parameters: This technique adds extra trainable parameters to LLMs, training them specifically on a dataset of edits while keeping the original parameters intact [45, 46]. (4) Adding Extra Modules: This strategy memorizes all editing instances and leverages a retrieval mechanism to fetch the most relevant pieces of information for each query [47, 25, 48, 49, 19].
@@ 2.3 Mixture-of-Experts (MoE)
Figure 2: Illustration of the LiveEdit framework. The upper part illustrates the editing process of LiveEdit. At time step $t$, the representation of an editing sample $(v_{e_{t}},p_{e_{t}},o_{e_{t}})$ at layer $l_{e}$ serves as an editing signal to generate the editing expert $(U_{e_{t}},V_{e_{t}})$ via $f_{eg}$ and routing features $(\hat{\phi}_{v_{e_{t}}},\hat{\psi}_{p_{e_{t}}})$ via $\hat{f}_{fe}$. Both are then added to the expert repository. The lower part shows the VLLM inference process with LiveEdit, where $\bar{f}_{fe}$ extracts input sample features at layer $l_{e}$ to route editing experts, which then adapt the representation.
@@ 4 The Proposed LiveEdit Framework
In this section, we formally introduce the LiveEdit framework, with the overall architecture shown in Figure 2. First, we explain how to generate corresponding experts for editing samples and how to extract the semantic features of the editing prompt. We then describe their use in extracting key visual features from the visual representation, thereby maintaining the expert repository. Next, we describe how LiveEdit routes experts during a single inference of the VLLM to instantly adjust its response. Finally, we elaborate on the overall training process of the LiveEdit model.
@@ 4.1 Construction and Update of Expert Repository
To perform both hard and soft routing for experts, we respectively extract semantics from the prompt representation and obtain key visual information from the visual representation through a feature extractor $\hat{f}_{fe}(\cdot)$:
@@ 4.2 Expert Routing and Editing on the Fly
Given $\mathcal{E}_{t}=\{(U_{e_{\tau}},V_{e_{\tau}},\hat{\phi}_{v_{e_{\tau}}},\hat{\psi}_{p_{e_{\tau}}})\}_{\tau=1}^{t}$, we filter for experts that are highly relevant to the input sample’s visual content by calculating the similarity between the key visual features of the input and the editing samples:
@@
where a trainable vision sentinel $\Theta\in\mathbb{R}^{N_{v}\times d}$ is set to dynamically determine the filtering threshold, following [20], which effectively avoids the bias caused by manually set thresholds. $N_{v}$ is the vision token count of $f_{\theta}$. Intuitively, if the input sample is more visually similar to the editing sample than the visual sentinel, then this editing sample should not be selected.
@@ 4.3 Training of LiveEdit
The training primarily consists of two parts: the editing loss, which ensures that the generated MoEs effectively guide the VLLM to follow the editing instructions, and the routing loss, which ensures hard and soft MoE routing. Given a batch of editing samples $\mathcal{D}_{e}=\{(v_{e_{b}},p_{e_{b}},o_{e_{b}})\}_{b=1}^{B}$, and their corresponding sampled generality and locality samples $\mathcal{D}_{g}=\{(v_{g_{b}},p_{g_{b}},o_{g_{b}})\}_{b=1}^{B}$ and $\mathcal{D}_{l}=\{(v_{l_{b}},p_{l_{b}},o_{l_{b}})\}_{b=1}^{B}$, the losses are formulated as follows.
4.3.1 Editing Loss
We mix the experts for the entire batch of editing samples to simulate the scenario during inference, where hard routing leads to multiple experts. First, we obtain the $l_{e}$-th layer outputs $\{(h_{v_{e_{b}}}^{l_{e}},h_{p_{e_{b}}}^{l_{e}},h_{o_{e_{b}}}^{l_{e}})\}_{b=1}^{B}$ for each part of an editing sample in $\mathcal{D}_{e}$. Then, through Eqs. 1 and 3, their corresponding experts and routing features can be obtained as $\{(U_{e_{b}},V_{e_{b}})\}_{b=1}^{B}$ and $\{(\hat{\phi}_{v_{e_{b}}},\hat{\psi}_{p_{e_{b}}})\}_{b=1}^{B}$. We define the expert set for soft routing fusion as ^=$\hat{\mathcal{E}}=\{(U_{e_{b}},V_{e_{b}},\hat{\psi}_{p_{e_{b}}})\}_{b=1}^{B}$. Thus, for any input sample, its representation at layer $l_{e}$ will be modified as in Eq. 6. Defining the VLLM modified in this way as $f_{\theta_{\hat{\mathcal{E}}}}$, the editing loss is defined as follows:
@@ 4.3.2 Routing Loss
In the routing part, we maximize the feature similarity between samples within the generality domain, while minimizing the feature similarity between unrelated samples. First, we randomly assign samples within the same generalization domain (i.e., editing samples and their corresponding generalization samples) into two new sets, defined as $\hat{\mathcal{D}}_{g}=\{[\mathcal{D}_{e}^{(b)},\mathcal{D}_{g}^{(b)}]_{\pi_{1}^{(b)}}\}_{b=1}^{B}$, and $\bar{\mathcal{D}}_{g}=\{[\mathcal{D}_{e}^{(b)},\mathcal{D}_{g}^{(b)}]_{\pi_{2}^{(b)}}\}_{b=1}^{B}$. Here, $\pi_{1}^{(b)},\pi_{2}^{(b)}\in[0,1]^{B}$ are the random integer vectors applied across the batch. This approach equalizes the reliability and generality of samples in feature matching with the edited and input samples, enhancing routing robustness. We use $\hat{f}_{fe}$ to extract the routing features $\{(\hat{\phi}_{g_{b}},\hat{\psi}_{g_{b}})\}_{b=1}^{B}$ from $\hat{\mathcal{D}}_{g}$ corresponding to edit end, and use $\bar{f}_{fe}$ to extract the routing features $\{(\bar{\phi}_{g_{b}},\bar{\psi}_{g_{b}})\}_{b=1}^{B}$ and $\{(\bar{\phi}_{l_{e}},\bar{\psi}_{l_{e}})\}_{b=1}^{B}$ from $\bar{\mathcal{D}}_{g}$ and $\mathcal{D}_{l}$ corresponding to input end, respectively. The routing loss is formulated as follows:
@@
Table 1: Partial results of lifelong edit performance for BLIP2-OPT and LLaVA-V1.5 on the E-VQA and VLKEB datasets. Due to space limitations, please refer to Appendix 8 for the complete results, including those for the E-IC dataset and the MiniGPT-4 model. “Rel.”, “T/M-Gen.” and “T/M-Loc.” stand for reliability, text/modal generality, and text/modal locality, respectively. “# Edit” indicates the number of edits. The t-tests demonstrate our improvements are statistically significant with $p<0.05$ level.
@@ 5.2 General Performance of Lifelong Editing
Table 1 shows partial results of lifelong editing experiments. In single-edit scenarios, our method generally achieves optimal performance. Methods such as FT-L, FT-M, and MEND [42] try to modify original model parameters and perform well initially. However, their performance deteriorates with more edits due to overfitting and cumulative parameter scaling [21, 58]. TP [45] addresses parameter scaling by introducing additional neurons, but a single neuron can’t encapsulate each edit’s visual information. Retrieval-based methods (SERAC [47], LTE [19], RECIPE [20]) remain robust in lifelong editing by decoupling edits from model parameters, yet struggle with semantically similar but visually different edit samples. LEMoE [46], a MoE-based editor, excels with few edits but suffers from issues like disruptive greedy routing and limited generalization due to overfitting experts to batch edits. Our method, LiveEdit, surpasses in edit performance. Contrastive learning-based expert routing resolves LEMoE’s issues, enhancing both edit speed and generalization. By decoupling edit samples into independent experts and applying a fusion strategy for vision-related experts, we prevent the semantic conflicts LEMoE encounters from its sequential batching. Importantly, our method maintains nearly 100% locality performance even with increasing edits. Three main factors contribute to our performance: First, hard routing filters out visually irrelevant experts. Second, soft routing scales influence by assigning lower scales to textually irrelevant experts. Finally, the locality editing loss further confines experts’ influence on response adaptation.
Figure 3: The impact of the combination of module dimension $d_{m}$ and expert rank $r$ on LiveEdit’s edit performance. Experiments are conducted on the E-VQA dataset, with 1,000 edits on BLIP2. Circle size represents LiveEdit’s training parameters, and color intensity indicates the average edit performance across five metrics.
@@ 5.4 Ablation Study
Figure 6: Instance analysis. The left side shows the model input to LLaVa after editing with LiveEdit. The right side displays the top 3 experts based on hard routing. The heatmaps represent the visual regions focused on when extracting key visual features. The bar chart below shows fusion weights from soft routing.

### v2

@@ Abstract
Model editing aims to correct inaccurate knowledge, update outdated information, and incorporate new data into Large Language Models (LLMs) without the need for retraining. This task poses challenges in lifelong scenarios where edits must be continuously applied for real-world applications. While some editors demonstrate strong robustness for lifelong editing in pure LLMs, Vision LLMs (VLLMs), which incorporate an additional vision modality, are not directly adaptable to existing LLM editors. In this paper, we propose LiveEdit, a Lifelong vision language model Edit to bridge the gap between lifelong LLM editing and VLLMs. We begin by training an editing expert generator to independently produce low-rank experts for each editing instance, with the goal of correcting the relevant responses of the VLLM. A hard filtering mechanism is developed to utilize visual semantic knowledge, thereby coarsely eliminating visually irrelevant experts for input queries during the inference stage of the post-edited model. Finally, to integrate visually relevant experts, we introduce a soft routing mechanism based on textual semantic relevance to achieve multi-expert fusion. For evaluation, we establish a benchmark for lifelong VLLM editing. Extensive experiments demonstrate that LiveEdit offers significant advantages in lifelong VLLM editing scenarios. Further experiments validate the rationality and effectiveness of each module design in LiveEdit.  [^1] The source code is available at https://github.com/qizhou000/LiveEdit.
@@ 1 Introduction
Generation of Low-Rank Experts: Mixture-of-Experts (MoE) combines multiple ”experts,” each specializing in specific data patterns or sub-tasks [26, 27]. In VLLM editing, we treat VLLM’s adherence to a single edit sample as a sub-task, where each new edit sample corresponds to a low-rank expert that adjusts the model’s response. MoE components are typically trained on sub-tasks, but directly fitting on individual edit samples results in poor generalization and inefficiency. To solve this, we propose an expert generator that creates low-rank experts for each new edit sample. The generator is trained to align VLLM with key edit metrics: reliability, textual/modal generality, and textual/modal locality [22]. The generated experts are stored in an expert repository. For new inputs, LiveEdit will select relevant experts from the repository and combines their adapted responses using hard and soft routing, as detailed below.
Hard and Soft Routing: We propose a two-stage routing strategy for expert utilization during inference. Attribution in [24] shows that VLLM processes prompts in early layers and extracts key visual features in later layers. In the first phase, we perform a text-to-vision interaction to extract key visual features and filter out noise. For each incoming sample, we compare its extracted features with those of edit samples, routing to visually relevant experts while filtering out visually irrelevant ones. Since this hard routing only considers visual semantics, it may select multiple visually matched but text-irrelevant experts. In the second phase, we apply soft routing through multi-expert fusion, incorporating the semantic similarity between the input query and the edit text. By combining absolute and relative weights, relevant edit samples are assigned higher weights and irrelevant ones lower weights. This approach suppresses text-irrelevant experts and avoids redundant interactions.
@@ 2.2 Model Editing
Model Editing for LLMs: We classify LLM editing into four categories. (1) Locate-Then-Edit methods identify and modify specific model parameters related to target knowledge [40]. ROME [17] uses causal mediation analysis for localization, while MEMIT [18] and WILKE [21] extend it for multi-editing. (2) Meta Learning methods employ a hyper-network to generate updated weights for edits  [41, 42, 43, 44]. (3) Additional Parameters methods introduce trainable parameters dedicated to edits while preserving original weights [45, 46]. Amon them, LEMOE [46] is based on MoE, but its greedy routing harms old experts’ influence when integrating new ones. (4) Adding Extra Modules methods store and retrieve edits via external memory mechanisms  [47, 25, 48, 49, 19]. In lifelong editing, the accumulation of shifts in the first two types hinders performance, while the latter two mitigate this by adding extra parameters and decoupling edits. However, in VLLM, extra modality and noise reduce efficacy.
@@ 2.3 Mixture-of-Experts (MoE)
Figure 2: Illustration of the LiveEdit framework. The upper part illustrates the editing process of LiveEdit. At time step $t$, the representation of an edit sample $(v_{e_{t}},p_{e_{t}},o_{e_{t}})$ at layer $l_{e}$ serves as an editing signal to generate the editing expert $(U_{e_{t}},V_{e_{t}})$ via $f_{eg}$ and routing features $(\hat{\phi}_{v_{e_{t}}},\hat{\psi}_{p_{e_{t}}})$ via $\hat{f}_{fe}$. Both are then added to the expert repository $\mathcal{E}_{t}$. The lower part shows the VLLM inference process with LiveEdit, where $\bar{f}_{fe}$ extracts input sample features at layer $l_{e}$ to route editing experts, which then adapt the representation.
@@ 4 The Proposed LiveEdit Framework
In this section, we formally introduce the LiveEdit framework, with the overall architecture shown in Figure 2. First, we explain how to generate corresponding experts for edit samples and how to extract the semantic features of the editing prompt. We then describe their use in extracting key visual features from the visual representation, thereby maintaining the expert repository. Next, we describe how LiveEdit routes experts during a single inference of the VLLM to instantly adjust its response. Finally, we elaborate on the overall training process of the LiveEdit model.
@@ 4.1 Construction and Update of Expert Repository
To perform both hard and soft routing for experts, we respectively extract key visual feature $\hat{\phi}_{v_{e_{t}}}$ using prompt semantics and the pure prompt feature $\hat{\psi}_{p_{e_{t}}}$ through a feature extractor $\hat{f}_{fe}(\cdot)$:
@@ 4.2 Expert Routing and Editing on the Fly
Given $\mathcal{E}_{t}=\{(U_{e_{\tau}},V_{e_{\tau}},\hat{\phi}_{v_{e_{\tau}}},\hat{\psi}_{p_{e_{\tau}}})\}_{\tau=1}^{t}$, we filter for experts that are highly relevant to the input sample’s visual content by calculating the similarity between the key visual features of the input and the edit samples:
@@
where a trainable vision sentinel $\Theta\in\mathbb{R}^{N_{v}\times d}$ is set to dynamically determine the filtering threshold, following [20], which effectively avoids the bias caused by manually set thresholds. $N_{v}$ is the vision token count of $f_{\theta}$. Intuitively, if the input sample is more visually similar to the edit sample than the visual sentinel, then this edit sample should not be selected.
@@ 4.3 Training of LiveEdit
The training primarily consists of two parts: the edit loss, which ensures that the generated MoEs effectively guide the VLLM to follow the editing instructions, and the routing loss, which ensures hard and soft MoE routing. Given a batch of edit samples $\mathcal{D}_{e}=\{(v_{e_{b}},p_{e_{b}},o_{e_{b}})\}_{b=1}^{B}$, and their corresponding sampled generality and locality samples $\mathcal{D}_{g}=\{(v_{g_{b}},p_{g_{b}},o_{g_{b}})\}_{b=1}^{B}$ and $\mathcal{D}_{l}=\{(v_{l_{b}},p_{l_{b}},o_{l_{b}})\}_{b=1}^{B}$, the losses are formulated as follows.
4.3.1 Edit Loss
We mix the experts for the entire batch of edit samples to simulate the scenario during inference, where hard routing leads to multiple experts. First, we obtain the $l_{e}$-th layer outputs $\{(h_{v_{e_{b}}}^{l_{e}},h_{p_{e_{b}}}^{l_{e}},h_{o_{e_{b}}}^{l_{e}})\}_{b=1}^{B}$ for each part of an edit sample in $\mathcal{D}_{e}$. Then, through Eqs. 1 and 3, their corresponding experts and routing features can be obtained as $\{(U_{e_{b}},V_{e_{b}})\}_{b=1}^{B}$ and $\{(\hat{\phi}_{v_{e_{b}}},\hat{\psi}_{p_{e_{b}}})\}_{b=1}^{B}$. We define the expert set for soft routing fusion as ^=$\hat{\mathcal{E}}=\{(U_{e_{b}},V_{e_{b}},\hat{\psi}_{p_{e_{b}}})\}_{b=1}^{B}$. Thus, for any input sample, its representation at layer $l_{e}$ will be modified as in Eq. 6. Defining the VLLM modified in this way as $f_{\theta_{\hat{\mathcal{E}}}}$, the edit loss is defined as follows:
@@ 4.3.2 Routing Loss
In the routing part, we maximize the feature similarity between samples within the generality domain, while minimizing the feature similarity between unrelated samples. First, we randomly assign samples within the same generalization domain (i.e., edit samples and their corresponding generality samples) into two new sets, defined as $\hat{\mathcal{D}}_{g}=\{[\mathcal{D}_{e}^{(b)},\mathcal{D}_{g}^{(b)}]_{\pi_{1}^{(b)}}\}_{b=1}^{B}$, and $\bar{\mathcal{D}}_{g}=\{[\mathcal{D}_{e}^{(b)},\mathcal{D}_{g}^{(b)}]_{\pi_{2}^{(b)}}\}_{b=1}^{B}$. Here, $\pi_{1}^{(b)},\pi_{2}^{(b)}\in[0,1]^{B}$ are the random integer vectors applied across the batch. This approach equalizes the reliability and generality of samples in feature matching with the edited and input samples, enhancing routing robustness. We use $\hat{f}_{fe}$ to extract the routing features $\{(\hat{\phi}_{g_{b}},\hat{\psi}_{g_{b}})\}_{b=1}^{B}$ from $\hat{\mathcal{D}}_{g}$ corresponding to edit end, and use $\bar{f}_{fe}$ to extract the routing features $\{(\bar{\phi}_{g_{b}},\bar{\psi}_{g_{b}})\}_{b=1}^{B}$ and $\{(\bar{\phi}_{l_{b}},\bar{\psi}_{l_{b}})\}_{b=1}^{B}$ from $\bar{\mathcal{D}}_{g}$ and $\mathcal{D}_{l}$ corresponding to input end, respectively. The routing loss is formulated as follows:
@@
Table 1: Partial results of lifelong edit performance for BLIP2-OPT and LLaVA-V1.5 on the E-VQA and VLKEB datasets. Due to space limitations, please refer to Appendix 8.1 for the complete results, including those for the E-IC dataset and the MiniGPT-4 model. “Rel.”, “T/M-Gen.” and “T/M-Loc.” stand for reliability, text/modal generality, and text/modal locality, respectively. “# Edit” indicates the number of edits. The t-tests demonstrate our improvements are statistically significant with $p<0.05$ level.
@@ 5.2 General Performance of Lifelong Editing
Table 1 shows partial results of lifelong editing experiments. In single-edit scenarios, our method generally achieves optimal performance. Methods such as FT-L, FT-M, and MEND [42] try to modify original model parameters and perform well initially. However, their performance deteriorates with more edits due to overfitting and cumulative parameter scaling [21, 58]. TP [45] addresses parameter scaling by introducing additional neurons, but a single neuron can’t encapsulate each edit’s visual information. Retrieval-based methods (SERAC [47], LTE [19], RECIPE [20]) remain robust in lifelong editing by decoupling edits from model parameters, yet struggle with semantically similar but visually different edit samples. LEMoE [46], a MoE-based editor, excels with few edits but suffers from issues like disruptive greedy routing and limited generalization due to overfitting experts to batch edits. Our method, LiveEdit, surpasses in edit performance. Contrastive learning-based expert routing resolves LEMoE’s issues, enhancing both edit speed and generalization. By decoupling edit samples into independent experts and applying a fusion strategy for vision-related experts, we prevent the semantic conflicts LEMoE encounters from its sequential batching. Importantly, our method maintains nearly 100% locality performance even with increasing edits. Three main factors contribute to our performance: First, hard routing filters out visually irrelevant experts. Second, soft routing scales influence by assigning lower scales to textually irrelevant experts. Finally, the locality edit loss further confines experts’ influence on response adaptation.
Figure 3: The impact of module dimension $d_{m}$ and expert rank $r$ on LiveEdit’s edit performance. Experiments are conducted on the E-VQA dataset, with 1,000 edits on BLIP2. Circle size represents LiveEdit’s training parameters, and color intensity indicates the average edit performance across five metrics.
@@ 5.4 Ablation Study
Figure 6: Instance analysis. The left side shows the model input to LLaVa after editing with LiveEdit. The right side displays the top 3 experts based on hard routing. The heatmaps represent the visual regions focused on when extracting key visual features. The bar chart below shows fusion weights from soft routing. Please refer to Appendix 8.2 for more analysis.
@@ 6 Conclusion
Acknowledgments. This work is supported by the National Key R&D Program of China (2022ZD0120302).
@@ 7.4 Model Settings and Training Details
8 More Experiments
8.1 The Complete Lifelong Editing Results
Figure 7: Instance analysis with difficult samples.
8.2 Instance Analysis with Difficult Samples
Figure 7 demonstrates the necessity and complementarity of hard and soft routing. Hard routing filters out irrelevant options based on key visuals, preventing the third-ranked expert from receiving excessive weight in soft routing. Then, soft routing adjusts by reducing the weight of the top-ranked expert that was overlooked by hard routing. Therefore, relying solely on the top expert selected by hard routing can lead to errors.
9 Other Discussions
Balance of Multi-Objective Losses: The challenge in multi-goal optimization lies in balancing conflicting objectives. As shown in Fig. 2, the Hard Routing Loss (HRL) doesn’t share parameters with Soft Routing/Editing Loss (SRL/EL), allowing it to be optimized separately. For EL and SRL, since EL can only be minimized by selecting the correct expert in a batch, and SRL helps EL choose the right expert, they align without conflict. In EL, we align with previous work [24]. In SRL, both losses boost semantic similarity. Consequently, we assign all loss weights as 1, and LiveEdit’s performance confirms the effectiveness of this choice.
Time and Space Overhead of LiveEdit: The edit time and additional memory required for each new expert remain constant, regardless of the sample size. For example, with LLaVA, it takes approximately 0.32 seconds and 160KB of memory on an A800 GPU.
Discussion with MoVA: The key differences between MoVA [71] and LiveEdit are as follows: 1) Task Focus: MoVA emphasizes macro-level task performance, whereas LiveEdit focuses on the micro-level control of specific responses. 2) Intervention: MoVA routes vision encoders externally to the LLM, while LiveEdit operates internally. 3) Routing: In the coarse stage, MoVA uses In-Context Learning (ICL) to filter out task-irrelevant vision experts, whereas LiveEdit removes vision-irrelevant edit experts. In the fine stage, MoVA reduces single-expert bias, while LiveEdit minimizes the influence of prompt-irrelevant edit experts, complementing the previous stage.